When you lift a ball with a force of 82 N, the ball accelerates upward with an acceleration of a. If you lift with a force of 92N, the ball's acceleration is 2a. Find the weight of the ball, and the acceleration a.


1
Expert's answer
2020-07-27T09:36:21-0400
"m(a)=F_1-mg\\m(2a)=F_2-mg\\m(a)=82-9.8m\\m(2a)=92-9.8m\\"

"m=7.3\ kg,a=1.4\frac{m}{s^2}"